40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jul 16 16:00:01 2019, gautam, Update, CDS, c1iscaux Supermicro setup IMG_7769.JPG
    Reply  Wed Jul 17 21:22:41 2019, gautam, Update, CDS, CM board Latch Enable subtlety 
       Reply  Sun Jul 21 12:55:38 2019, gautam, Update, CDS, CM board Latch Enable test script LatchLogic.pdfLatchLogicTest.png
          Reply  Fri Aug 9 08:59:04 2019, gautam, Update, CDS, Prep for install of c1iscaux 
             Reply  Sun Aug 11 11:47:42 2019, gautam, Update, CDS, Bench test of c1iscaux 
                Reply  Mon Aug 12 17:36:04 2019, gautam, Update, CDS, More bench test of c1iscaux iscauxCheclist.pdf
                   Reply  Mon Aug 12 21:25:19 2019, Koji, Update, CDS, More bench test of c1iscaux 
                   Reply  Tue Aug 13 08:07:09 2019, gautam, Update, CDS, P1--->P2 
                      Reply  Tue Aug 13 14:36:17 2019, gautam, Update, CDS, P1--->P2 
                   Reply  Fri Aug 16 16:40:04 2019, gautam, Update, CDS, 1Y3 work newLook1Y3.JPGIMG_7803.JPGc1lsc_crashed.png
                      Reply  Sat Aug 17 16:49:23 2019, gautam, Update, CDS, More 1Y3 work Screen_Shot_2019-08-17_at_3.00.57_PM.pngScreen_Shot_2019-08-17_at_3.12.23_PM.pngIMG_7804.JPGScreenshot_from_2019-08-17_17-04-47.png
                         Reply  Mon Aug 19 14:36:21 2019, gautam, Update, CDS, c1iscaux remaining work caseForSmallerFootprint.pdf
                            Reply  Tue Aug 20 19:05:24 2019, Koji, Update, CDS, MC1 (and MC3) troubleshoot Screenshot_from_2019-08-20_17-26-01.pngScreenshot_from_2019-08-20_17-43-03.png
                               Reply  Thu Aug 22 12:54:06 2019, Koji, Update, CDS, MC1 glitch removed (for now) and IMC locking recovered 7x
                                  Reply  Thu Aug 22 20:56:51 2019, Koji, Update, CDS, MC1 glitch removed (for now) and IMC locking recovered 
                            Reply  Fri Aug 23 18:46:17 2019, Jon, Update, CDS, c1iscaux remaining work 
                               Reply  Sun Aug 25 14:18:08 2019, gautam, Update, CDS, c1iscaux remaining work Screen_Shot_2019-08-25_at_10.38.37_PM.png
    Reply  Thu Jul 18 10:46:04 2019, gautam, Update, CDS, Database files made 
       Reply  Thu Jul 18 19:58:56 2019, gautam, Update, CDS, Work on Acromag chassis IMG_7771.JPGIMG_7770.JPGIMG_7772.JPG
          Reply  Fri Jul 19 19:44:03 2019, gautam, Update, CDS, Database file test Whitening.png
Message ID: 14857     Entry time: Sun Aug 25 14:18:08 2019     In reply to: 14855
Author: gautam 
Type: Update 
Category: CDS 
Subject: c1iscaux remaining work 

There were a bunch of useless / degenerate channels added - e.g. whitening gains which are alreay burt-snapshot. Maybe there are many more useless channels being trended, but no need to add more.

Copy-pasting wasn't done correctly - the first 4 added channels were duplicates. There are in fact 5 LO power mons, one for each of the frequencies 11, 33, 55, 110 and 165 MHz. 

I cleaned up. Basically only the detect-mon channels, and the ALS channels, are new in the setup now. I will review if any extra channels are required later. While checking that the daqd is happy, I noticed c1lsc FEs are in their stuck state, see Attachment #1. I guess a cable was bumped when the strain relief operation was underway. I'm not attempting a remote resuscitation.

Quote:

I added the list of new c1iscaux channels to /opt/rtcds/caltech/c1/chans/daq/C0EDCU.ini and restarted the framebuilder. Koji had thought some of these channels might have previously existed under slightly different names. However, after looking through C0EDCU.ini and the other _SLOW.ini files, I did not find any candidates for removal. As far as I can tell, all of these channels are being recorded for the first time.

Attachment 1: Screen_Shot_2019-08-25_at_10.38.37_PM.png  206 kB  Uploaded Sun Aug 25 15:39:52 2019  | Hide | Hide all
Screen_Shot_2019-08-25_at_10.38.37_PM.png
ELOG V3.1.3-